calculate pH obtained by mixing equal volumes of 0.1M h2so4 and 0.3M Naoh


1
Expert's answer
2021-12-20T17:06:03-0500

First start by writing the correct equation.

"2NaOH+H_2SO_4\longrightarrow Na_2SO_4+2H_2O"

From the equation, 2 moles of sodium hydroxide reacts with a mole of sulphuric acid.

"2\space moles\space NaOH \equiv 1\space mole\space sulphuric\space acid"

Using the above relation we can find the moles of sodium hydroxide that will react as

"\frac{0.1×2}{1}=0.2\space moles."

The moles of sodium hydroxide are in excess and the remaining amount is

"0.3-0.2=0.1\space moles"

The pH of the resulting solution is

"pOH=(-log\space 0.1)=1\\pH=14-1=13\\pH=13"